OpenStudy (anonymous):

limits question, please help! equation will be be posted in a few minutes on the equation editor.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \left[ (2+h)^3-8 \right]\div h\]

OpenStudy (anonymous):

have you learned l'hospitals rule yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

I don't see the limit problem

OpenStudy (anonymous):

thats not hopitals rule...

OpenStudy (anonymous):

its right under your first reply.

OpenStudy (anonymous):

basically l'hospitals rule says that if the limit is indeterminant and is in a form F(x)/G(x) then the limit can be found by taking F'(x)/G'(x)

OpenStudy (anonymous):

you need to cube it out: (2+h)(2+h)(2+h) = h^3 + 6h^2 + 12h +8 So now when we subtract 8, we are left with all terms that have an h attached to them. So cancel an h from every term and get rid of the h in the denominator and then plug in h=0 to get a final answer of: 12

OpenStudy (anonymous):

thank you! :)